4x^3+3x^2=64x+48
I tried and ended up with x=16,-16, and -3/4. Is that correct?

Answer by Alan3354(69443)

4x^3+3x^2=64x+48
I tried and ended up with x=16,-16, and -3/4. Is that correct?
----------
Sub 16 for x and see.
4*16^3 + 3*16^2 =? 64*16 + 48
16384 + 768 =? 1024 + 48
No.